Suppose that a firm produces 200,000 units a year and sells them all for $10 each. The explicit costs of production are $1,500,000 and the implicit costs of production are $300,000. The firm has an accounting profit of:

a)     $500,000 and an economic profit of $200,000

b)     $400,000 and an economic profit of $200,000

c)     $300,000 and an economic profit of $400,000

d)     $200,000 and an economic profit of $500,000

 


1
Expert's answer
2021-04-19T07:34:49-0400

Accounting Profit= Total Revenue-Explicit Cost

Total Revenue= 200,000*10= 2,000,000

Explicit Cost= 1,500,00

Thus, 2,000,000-1,500,000= 500,000

Economic Profit= Total Revenue- (Explicit+Implicit Cost)

Implicit Cost= 300,000

Thus, 2,000,000-(1,500,000+300,000)=200,000

The Answer is A